I have tried this problem about 10 different ways but I can't seem to get it!! Help!!!
A merchant has coffee worth $20 a pound that she wishes to mix with 30 pounds of coffee worth $90 a pound to get a mixture that can be sold for $40 a pound. How many pounds of the $20 coffee should be used?

Answer by edjones(7311)

Let x be the pounds of $20 coffee necessary.
.
30*90=2700
x *20=20x
(30+x)*40=1200+40x
.
20x+2700=40x+1200
20x=1500
x=75 lbs.
.
Check:
75+30=105 lbs
75*20=1500
1500+2700=4200
105*40=4200
.
Ed